the mean and SD of 6 observations are 8 and 4 respectively if it has a observation is multiplied by by 3 then the new mean and SD of the resulting observations

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
when multiplied each observation by a constant a , mean will be multiplied by that constant.
 
hence new mean = 8* 3 = 24
 
similar case with standard deviation
 
new standard deviation = 3 * 4 = 12